OpenStudy (anonymous):

If R, S, and Q can wallpaper a house in 8 hours and R and S can do it in 12 hours, how long will it take Q alone to wallpaper the house?

OpenStudy (anonymous):

T=(ab)/(a+b) 8=(12Q)/(12+Q) 8(12+Q)=12Q 96+8Q=12Q 96=4Q Q=24 hours